OpenStudy (skullcandy229):

Find the difference between -14w - 3 and 5w I will give you a medal

OpenStudy (anonymous):

The numerical meaning for "difference" is "subtraction." Ex: The difference between 5 and 3 is 2 Mathematically: 5-3=2

OpenStudy (anonymous):

I mean, the mathematical meaning*

OpenStudy (anonymous):

(-14w - 3) - (5w) = (-14w - 5w) - 3 = -19w -3

OpenStudy (skullcandy229):

o

OpenStudy (anonymous):

@livias.random Is correct, but I think you should also understand what the statement you're given is saying.

OpenStudy (anonymous):

@brooke2353 That's already been said.

OpenStudy (brooke2353):

OH MY BAD

OpenStudy (brooke2353):

I will delete it then

OpenStudy (anonymous):

@CShrix, don't worry, I understand, and I figured that since you did a great job of defining 'difference' mathematically repeating was unnecessary

OpenStudy (skullcandy229):

I whish I could give you all a medal